# Simulating Disky Broad Line Region Reverberation

### v1

@@ 2.4 Models for the Light Curve
We utilize light curve templates or synthetic light curves from statistical models in order to simulate reverberation of the BLR. The origin and uses of light curve templates is further discussed in Section 4.1. For the purposes of creating synthetic light curves, we utilize the statistical and empirical Damped Random Walk (DRW) model for light curves of AGNs (e.g., kelly09; macleod12). macleod10 and suberlak21 show that the model parameters of the DRW can be linked to physical properties of the central engine such as the black hole mass and luminosity of the AGN, and more details are given in Section 4.1. The use of the connections found by macleod10 and suberlak21 is vital in ensuring that the synthetic light curves created also remain physically realistic. Examples of light curves are shown in Section 4.1. The use of synthetic light curves allows us to create long, uninterrupted time series of output parameters and to explore the behavior of the models over a wide range of input parameters.
3 Overview of the Pipeline
@@
The pipeline assumes a specific geometry and specific emission properties for a disk-like BLR and takes in a light curve for the illuminating source. It then creates a time sequence of line profiles and subsequently measures model-independent shape parameters of these profiles. The process is illustrated in Figure 4. The light curve generator, discussed in Section 4.1, allows the user to choose a synthetic or empirical light curve template, and the emissivity modules discussed in Section 2.3 allow a choice of components for the final emissivity law. The reverberation module, discussed in Section 4.2, uses the adopted disk geometry and generates instantaneous apparent illumination maps using the corresponding geometrical time delay of the light as it crosses the BLR. The line profile module of Section 4.3 uses these brightness maps, as well as the disk geometry, and generates a time series of smooth emission line profiles. The profile parameter module in Section 4.4 measures shape parameters of each emission line profile. There is an option to add simulated noise to the smooth emission line profiles, which is discussed in Section 4.5. The end result is a time series of synthetic profiles and profile parameters that allows us to quantify how these broad emission lines change over time, and how they may ‘jitter’ or ‘breathe’.
4 Modules of the Pipeline
4.1 Light Curves
@@
As discussed in Section 2.4, a synthetic light curve is created using the statistical and empirical DRW model, and is part of the Statistical Model portion of Figure 4. However, in the future, other models for AGN light curves may also be utilized in this module. There are three input parameters for the DRW model: the structure function at infinity , a characteristic timescale , and a mean absolute magnitude in the i-band , as described in kelly09 and macleod10. This model successfully recreates the photometric variability of AGNs and nearby quasars. macleod10 and suberlak21 found that there is no dependence of the input parameters on redshift, but there are empirical relations between the rest-frame wavelength, black hole mass, and absolute mean-magnitude in the i-band. The damped random walk parameters can be written as
@@
where represents the parameter of interest, or . The coefficients , , , take different values depending on which damped random walk parameter we wish to compute. We utilize the prescriptions in kelly09 for implementing this auto-regressive process.
@@
In order to get an absolute i-band magnitude, we calculate the bolometric luminosity of the AGN under the assumption that its black hole is accreting with a chosen Eddington ratio, . We use a correction factor described in richards06 to get the i-band luminosity, with a median wavelength of 7625 Å for the i-band. We chose a rest-frame wavelength of approximately 912 Å to model the ionizing continuum that is primarily responsible for the line emission. This is vital in ensuring that the light curve generated refers to the energy band just above 1 Ry. Examples of light curves with varying input parameters can be seen in Figure 5. Testing of the parameter space of the model will utilize these general synthetic light curves.
@@
Empirical light curve templates can also be used in the pipeline for testing specific scenarios for comparison to observed parameters. It is important to note that empirical templates undergo additional processing in the Light Curve Generator Module seen in Figure 4. The pipeline ‘stretches’ the amplitude of the fluctuations as studies (e.g., macleod12; sun15) have shown that at shorter wavelengths, the variability amplitude increases, a trend known commonly as ‘bluer when brighter’. Vertical stretching of the light curve, therefore, gives a better representation of the ionizing continuum, and produces the same shape in the light curve but with more pronounced fluctuations, as demonstrated in Figure 6.
@@
An empirical light curve used for testing against observations was the measured continuum of NGC 5548 obtained in an intensive reverberation mapping campaign, as detailed in kriss19 and shown in Figure 6. The light curve was produced from flux observations at 5100 Å and discussed in derosa18.
4.2 Reverberation
@@
Under the assumption that the BLR is a flat disk centered on the continuum source, we use the formula from peterson93 for the time delay:
@@
We calculate the inner and outer radii of the BLR based on physical considerations. Using the empirical relation described in bentz13, the characteristic radius is related to the luminosity through
@@
where and are the best fit values from the study and is the luminosity measured at 5100 Å. richards06 provide a conversion factor, which was , to connect bolometric luminosity to the luminosity measured at 5100 Å, which is used in equation (13). The outer radius is set to be five times this value as studies such as koshida14, yang20, and lira24 have found that dust reverberation from the torus has time delays that are approximately five times longer than the BLR reverberation timescale. The inner radius of the BLR for the tests carried out here is set to be exactly 500 , which is another empirical relation found in storchi-bergmann17.
@@
In the Reverberation module in Figure 4, we calculate a time delay for each location on the disk and connect that delay with a point on the light curve described in Section 4.1. The module assigns a brightness or instantaneous illumination for a particular location in the BLR by using the corresponding intensity of the light curve and the light travel time to the BLR pixel. The result is a reverberation map of the disk for one particular instant, with an example shown in Figure 7. This, and other reverberation maps, are proportional to the flux arriving at a specific BLR element.
@@
The reverberation map is multiplied by the apparent emissivity map discussed in Section 2.3 to produce the instantaneous apparent brightness map. An example of this final map is shown in Figure 8. The instantaneous apparent brightness map is used in the calculation of the synthetic broad emission line profiles.
4.3 Line Profiles
@@
The instantaneous apparent brightness maps are passed to the Line Profile Module seen in Figure 4, whose original numerical methods and model were described in Section 2.2. Other input parameters include the inner and outer BLR radius, inclination angle, and rest wavelength of the emission line. The numerical integration grid uses logarithmic steps in radius to capture the rapid radial variation of emissivity and linear steps in azimuth. The output is a smooth, normalized emission line profile, as demonstrated in Figure 9. A separate profile is computed at each time step, corresponding to each instantaneous apparent brightness map. For the purposes of setting up and testing the pipeline, we have simulated the broad H line, as seen in Figure 9.
4.4 Profile Parameter Module
@@
One can measure shape parameters from the emission line profiles in order to characterize them using the Profile Parameter Module in Figure 4. These parameters can be extracted before and after the addition of the noise in the profiles in order to examine the systematic distortions produced by noise (discussed in Section 4.5). There are three parameters of interest, which are associated with the first three moments of the profile: the centroid velocity shift, the velocity dispersion, and the Pearson skewness coefficient.
@@
The Profile Parameter Module is the last module in the pipeline, and the user is left with multiple outputs. As seen in Figure 4, there are the time series of instantaneous brightness maps, the time series of the ideal line profiles, the time series of the noisy profiles (if choosing to add noise), and the time series of synthetic profile parameters which can be further used for other tests. An illustrative example of these parameters and a discussion of a full test of the pipeline can be found in Section 5.2.
4.5 Synthetic Noise
@@
Here, is a normalization constant and is the fiducial wavelength at which the user specifies the signal-to-noise ratio (S/N). The power-law index can range from to 3 (edelson86). For the purposes of the experiments below, we chose and normalize at 6563 Å. The relative strength of the continuum is an adjustable parameter and the S/N scales with the square root of the flux density to mimic Poisson noise. We obtain the noisy value through the use of a random normal distribution generator. This normal distribution has a mean equal the local flux at any given pixel in the spectrum, and a standard deviation calculated using the local value of the S/N. An example of a noisy emission line profile is shown in Figure 10.
@@
5 Discussion and Conclusions
5.1 Limitations of the Pipeline
The current limitations of the pipeline are set by the assumptions listed in Section 2. Some of the limiting assumptions are that the disk is thin, and that the spiral arm is a brightness perturbation. The use of spiral arms is a departure from the assumption that the gas moves in a simple circular and Keplerian velocity field, and the departures from Keplerian rotation associated with the spiral arm are not included.
@@ 5.1 Limitations of the Pipeline
where is the line emissivity and is the flux of the ionizing continuum at some fiducial wavelength impinging on the surface of the disk. We assume that the responsivity is linear everywhere, i.e., a given fractional change in the intensity of the ionizing continuum produces the same fractional change in the emissivity of the line at that location. Future work on this project will incorporate a non-linear responsivity of the disk using methods described in korista04 and goad14. The use of non-linear responsivities would reflect more accurately the observations of the mean and rms spectra from reverberation mapping campaigns, as discussed in li25.
5.2 Tests of the Pipeline
@@
We have carried out two end-to-end tests of the pipeline using both a synthetic light curve and an empirical template. Both tests have a S/N of 40 for the continuum, and the rms spread is calculated as described in Section 4.5. More specific information regarding the values for each test can be found in the captions for Figures 12 and 13. The first test utilizes the light curve template presented in Figure 6, and has the same disk parameters as those presented in Figure 1, which makes the disk axisymmetric. This test produces the time series of profile parameters presented in Figure 12. Velocity jitter, as discussed in Section 1 and seen in Figure 12b, represents the fluctuations of the centroid of the line profile with time. We see that all three parameters rise sharply at the beginning in response to driving/ionizing light curve. Furthermore, we find that the line profile parameter time series seem to resemble the input driving light curve with only a small time delay-this reaffirms the process of our ‘reverberation’ and may illustrate that the underlying emissivities and Keplerian velocity distribution are not the dominant process in the changes in the emission line profiles. Studies such as barth15 have found that velocity jitter can appear on BLR light-crossing timescales (days to weeks), and that the pattern of these changes seems to be correlated with the shape of the continuum light curve, which is presented in Figure 12a. After a visual examination of the centroid velocity shifts shown in Figure 12b, we confirm that the jitter has the same variation of the continuum light curve, which was seen in barth15. Similarly, the amplitudes of the centroid velocity shift are of the same magnitude as those observed by barth15. This test illustrates the current capabilities of the pipeline and the plausibility of the initial results. We also see that all of the velocity shifts are positive, which indicates that all of the profiles have been redshifted as a result of special and general relativistic effects incorporated in the line profile calculations. The observed centroid shift corresponds to the combined effect of the gravitational and transverse redshift at a radius of 5000 in the disk, which is near the middle of the assumed line-emitting annulus.
@@
5.3 Summary and Future Work
In this paper, we have presented a pipeline that can simulate reverberation from a disk-like BLR. We use physical models and analytic prescriptions to specify features of the BLR, such as the geometry, kinematics, and emissivity. We incorporate special and general relativistic effects in the calculation of the line profiles. We focus specifically on the disk-like geometry for the BLR, but this pipeline can be adapted to use other geometries and velocity fields as well. We produce synthetic broad emission line profiles and measure their intrinsic shape parameters to see the changes they exhibit over the light-crossing time of the BLR, which may help further inform us on the BLR kinematics and geometry. Our pipeline does not require significant computational power or computational time when running one realization of the pipeline, and averages approximately 8 minutes for one realization. This speed can help identify narrow parameter spaces that can be further investigated using other sophisticated forward-modeling codes introduced in Section 1.
In order to increase the speed of this pipeline, work is currently underway to parallelize some of the algorithms. Further work will explore the full parameter space of this model in order to investigate conditions that can cause the breathing or anti-breathing behavior, as well as the velocity jitter. Our forthcoming parameter space exploration will be guided by the range in disk parameters found by previous studies (see Section 2.3). As mentioned in Section 5.1, we will also incorporate a non-linear responsivity of the disk. Future applications will use long duration synthetic light curves that can span the dynamical time of the BLR. Using other empirical light curves, such as those from the Sloan Digital Sky Survey (SDSS) V Reverberation Mapping Campaign, can provide direct comparisons between the behaviors of our synthetic broad emission line profiles and those that are observed, while also allowing exploration of the model’s parameter space. With light curves expected from future telescopes and surveys such as the Long Synoptic Survey Telescope (ivezic19), we can utilize light curves with baselines as small as 200 days up to months to interpret observed line profiles. In a forthcoming paper, we will present the results of a parameter space exploration, as well as the consequences of different combinations of parameters for velocity jitter and breathing.

### v2

@@ II.4 Models for the Light Curve
We utilize light curve templates or synthetic light curves from statistical models in order to simulate reverberation of the BLR. The origin and uses of light curve templates is further discussed in Section IV.1. For the purposes of creating synthetic light curves, we utilize the statistical and empirical Damped Random Walk (DRW) model for light curves of AGNs (e.g., Kelly et al., 2009; MacLeod et al., 2012). MacLeod et al. (2010) and Suberlak et al. (2021) show that the model parameters of the DRW can be linked to physical properties of the central engine such as the black hole mass and luminosity of the AGN, and more details are given in Section IV.1. The use of the connections found by MacLeod et al. (2010) and Suberlak et al. (2021) is vital in ensuring that the synthetic light curves created also remain physically realistic. Examples of light curves are shown in Section IV.1. The use of synthetic light curves allows us to create long, uninterrupted time series of output parameters and to explore the behavior of the models over a wide range of input parameters.
III Overview of the Pipeline
@@
The pipeline assumes a specific geometry and specific emission properties for a disk-like BLR and takes in a light curve for the illuminating source. It then creates a time sequence of line profiles and subsequently measures model-independent shape parameters of these profiles. The process is illustrated in Figure 4. The light curve generator, discussed in Section IV.1, allows the user to choose a synthetic or empirical light curve template, and the emissivity modules discussed in Section II.3 allow a choice of components for the final emissivity law. The reverberation module, discussed in Section IV.2, uses the adopted disk geometry and generates instantaneous apparent illumination maps using the corresponding geometrical time delay of the light as it crosses the BLR. The line profile module of Section IV.3 uses these brightness maps, as well as the disk geometry, and generates a time series of smooth emission line profiles. The profile parameter module in Section IV.4 measures shape parameters of each emission line profile. There is an option to add simulated noise to the smooth emission line profiles, which is discussed in Section IV.5. The end result is a time series of synthetic profiles and profile parameters that allows us to quantify how these broad emission lines change over time, and how they may ‘jitter’ or ‘breathe’.
IV Modules of the Pipeline
IV.1 Light Curves
@@
As discussed in Section II.4, a synthetic light curve is created using the statistical and empirical DRW model, and is part of the Statistical Model portion of Figure 4. However, in the future, other models for AGN light curves may also be utilized in this module. There are three input parameters for the DRW model: the structure function at infinity , a characteristic timescale , and a mean absolute magnitude in the i-band , as described in Kelly et al. (2009) and MacLeod et al. (2010). This model successfully recreates the photometric variability of AGNs and nearby quasars. MacLeod et al. (2010) and Suberlak et al. (2021) found that there is no dependence of the input parameters on redshift, but there are empirical relations between the rest-frame wavelength, black hole mass, and absolute mean-magnitude in the i-band. The damped random walk parameters can be written as
@@
where represents the parameter of interest, or . The coefficients , , , take different values depending on which damped random walk parameter we wish to compute. We utilize the prescriptions in Kelly et al. (2009) for implementing this auto-regressive process.
@@
In order to get an absolute i-band magnitude, we calculate the bolometric luminosity of the AGN under the assumption that its black hole is accreting with a chosen Eddington ratio, . We use a correction factor described in Richards et al. (2006) to get the i-band luminosity, with a median wavelength of 7625 Å for the i-band. We chose a rest-frame wavelength of approximately 912 Å to model the ionizing continuum that is primarily responsible for the line emission. This is vital in ensuring that the light curve generated refers to the energy band just above 1 Ry. Examples of light curves with varying input parameters can be seen in Figure 5. Testing of the parameter space of the model will utilize these general synthetic light curves.
@@
Empirical light curve templates can also be used in the pipeline for testing specific scenarios for comparison to observed parameters. It is important to note that empirical templates undergo additional processing in the Light Curve Generator Module seen in Figure 4. The pipeline ‘stretches’ the amplitude of the fluctuations as studies (e.g., MacLeod et al., 2012; Sun et al., 2015) have shown that at shorter wavelengths, the variability amplitude increases, a trend known commonly as ‘bluer when brighter’. Vertical stretching of the light curve, therefore, gives a better representation of the ionizing continuum, and produces the same shape in the light curve but with more pronounced fluctuations, as demonstrated in Figure 6.
@@
An empirical light curve used for testing against observations was the measured continuum of NGC 5548 obtained in an intensive reverberation mapping campaign, as detailed in Kriss et al. (2019) and shown in Figure 6. The light curve was produced from flux observations at 5100 Å and discussed in De Rosa et al. (2018).
IV.2 Reverberation
@@
Under the assumption that the BLR is a flat disk centered on the continuum source, we use the formula from Peterson (1993) for the time delay:
@@
We calculate the inner and outer radii of the BLR based on physical considerations. Using the empirical relation described in Bentz et al. (2013), the characteristic radius is related to the luminosity through
@@
where and are the best fit values from the study and is the luminosity measured at 5100 Å. Richards et al. (2006) provide a conversion factor, which was , to connect bolometric luminosity to the luminosity measured at 5100 Å, which is used in equation (13). The outer radius is set to be five times this value for the purposes of testing the pipeline, as studies such as Koshida et al. (2014), Yang et al. (2020), and Lira et al. (2024) have found that dust reverberation from the torus has time delays that are approximately five times longer than the BLR reverberation timescale. The inner radius of the BLR for the tests carried out here is set to be exactly 500 , which is another empirical relation found in Storchi-Bergmann et al. (2017). The assumption we have made, especially the linear responsivity and monotonically decreasing emissivity, and the fixed values that we have adopted for the inner and outer radii are suitable only for testing the pipeline and do not allow us to self-consistently compute a “characteristic radius” of the BLR. We defer this exercise to future work.
@@
In the Reverberation module in Figure 4, we calculate a time delay for each location on the disk and connect that delay with a point on the light curve described in Section IV.1. The module assigns a brightness or instantaneous illumination for a particular location in the BLR by using the corresponding intensity of the light curve and the light travel time to the BLR pixel. The result is a reverberation map of the disk for one particular instant, with an example shown in Figure 7. This, and other reverberation maps, are proportional to the flux arriving at a specific BLR element.
@@
The reverberation map is multiplied by the apparent emissivity map discussed in Section II.3 to produce the instantaneous apparent brightness map. An example of this final map is shown in Figure 8. The instantaneous apparent brightness map is used in the calculation of the synthetic broad emission line profiles.
IV.3 Line Profiles
@@
The instantaneous apparent brightness maps are passed to the Line Profile Module seen in Figure 4, whose original numerical methods and model were described in Section II.2. Other input parameters include the inner and outer BLR radius, inclination angle, and rest wavelength of the emission line. The numerical integration grid uses logarithmic steps in radius to capture the rapid radial variation of emissivity and linear steps in azimuth. The output is a smooth, normalized emission line profile, as demonstrated in Figure 9. A separate profile is computed at each time step, corresponding to each instantaneous apparent brightness map. For the purposes of setting up and testing the pipeline, we have simulated the broad H line, as seen in Figure 9.
IV.4 Profile Parameter Module
@@
One can measure shape parameters from the emission line profiles in order to characterize them using the Profile Parameter Module in Figure 4. These parameters can be extracted before and after the addition of the noise in the profiles in order to examine the systematic distortions produced by noise (discussed in Section IV.5). There are three parameters of interest, which are associated with the first three moments of the profile: the centroid velocity shift, the velocity dispersion, and the Pearson skewness coefficient.
@@
The Profile Parameter Module is the last module in the pipeline, and the user is left with multiple outputs. As seen in Figure 4, there are the time series of instantaneous brightness maps, the time series of the ideal line profiles, the time series of the noisy profiles (if choosing to add noise), and the time series of synthetic profile parameters which can be further used for other tests. An illustrative example of these parameters and a discussion of a full test of the pipeline can be found in Section V.2.
IV.5 Synthetic Noise
@@
Here, is a normalization constant and is the fiducial wavelength at which the user specifies the signal-to-noise ratio (S/N). The power-law index can range from to 3 (Edelson and Malkan, 1986). For the purposes of the experiments below, we chose and normalize at 6563 Å. The relative strength of the continuum is an adjustable parameter and the S/N scales with the square root of the flux density to mimic Poisson noise. We obtain the noisy value through the use of a random normal distribution generator. This normal distribution has a mean equal the local flux at any given pixel in the spectrum, and a standard deviation calculated using the local value of the S/N. An example of a noisy emission line profile is shown in Figure 10.
@@
V Discussion and Conclusions
V.1 Limitations of the Pipeline
The current limitations of the pipeline are set by the assumptions listed in Section II. Some of the limiting assumptions are that the disk is thin, and that the spiral arm is a brightness perturbation. The use of spiral arms is a departure from the assumption that the gas moves in a simple circular and Keplerian velocity field, and the departures from Keplerian rotation associated with the spiral arm are not included.
@@ V.1 Limitations of the Pipeline
where is the line emissivity and is the flux of the ionizing continuum at some fiducial wavelength impinging on the surface of the disk. We assume that the responsivity is linear everywhere, i.e., a given fractional change in the intensity of the ionizing continuum produces the same fractional change in the emissivity of the line at that location. Future work on this project will incorporate a non-linear responsivity of the disk using methods described in Korista and Goad (2004) and Goad and Korista (2014). The use of non-linear responsivities would reflect more accurately the observations of the mean and rms spectra from reverberation mapping campaigns, as discussed in Li and Wang (2025).
V.2 Tests of the Pipeline
@@
We have carried out two end-to-end tests of the pipeline using both a synthetic light curve and an empirical template. Both tests have a S/N of 40 for the continuum, and the rms spread is calculated as described in Section IV.5. More specific information regarding the values for each test can be found in the captions for Figures 12 and 13. The parameters used for these tests were chosen to showcase the capabilities of the pipeline; as such they are generally plausible but do not represent any specific system. The first test utilizes the light curve template presented in Figure 6, and has the same disk parameters as those presented in Figure 1, which makes the disk axisymmetric. This test produces the time series of profile parameters presented in Figure 12. Velocity jitter, as discussed in Section I and seen in Figure 12b, represents the fluctuations of the centroid of the line profile with time. We see that all three parameters rise sharply at the beginning in response to driving/ionizing light curve. Furthermore, we find that the line profile parameter time series seem to resemble the input driving light curve with only a small time delay-this reaffirms the process of our ‘reverberation’ and may illustrate that the underlying emissivities and Keplerian velocity distribution are not the dominant process in the changes in the emission line profiles. Studies such as Barth et al. (2015) have found that velocity jitter can appear on BLR light-crossing timescales (days to weeks), and that the pattern of these changes seems to be correlated with the shape of the continuum light curve, which is presented in Figure 12a. After a visual examination of the centroid velocity shifts shown in Figure 12b, we confirm that the jitter has the same variation of the continuum light curve, which was seen in Barth et al. (2015). Similarly, the amplitudes of the centroid velocity shift are of the same magnitude as those observed by Barth et al. (2015). This test illustrates the current capabilities of the pipeline and the plausibility of the initial results. We also see that all of the velocity shifts are positive, which indicates that all of the profiles have been redshifted as a result of special and general relativistic effects incorporated in the line profile calculations. The observed centroid shift corresponds to the combined effect of the gravitational and transverse redshift at a radius of 5000 in the disk, which is near the middle of the assumed line-emitting annulus.
@@
V.3 Summary and Future Work
In this paper, we have presented a pipeline that can simulate reverberation from a disk-like BLR. We use physical models and analytic prescriptions to specify features of the BLR, such as the geometry, kinematics, and emissivity. We incorporate special and general relativistic effects in the calculation of the line profiles. We focus specifically on the disk-like geometry for the BLR, but this pipeline can be adapted to use other geometries and velocity fields as well. We produce synthetic broad emission line profiles and measure their intrinsic shape parameters to see the changes they exhibit over the light-crossing time of the BLR, which may help further inform us on the BLR kinematics and geometry. Our pipeline does not require significant computational power or computational time when running one realization of the pipeline, and averages approximately less than a minute for one realization, including the saving of intermediary steps. This speed can help identify narrow parameter spaces that can be further investigated using other sophisticated forward-modeling codes introduced in Section I.
In order to increase the speed of this pipeline, work is currently underway to parallelize some of the algorithms. Further work will explore the full parameter space of this model in order to investigate conditions that can cause the breathing or anti-breathing behavior, as well as the velocity jitter. Our forthcoming parameter space exploration will be guided by the range in disk parameters found by previous studies (see Section II.3). As mentioned in Section V.1, we will also incorporate a non-linear responsivity of the disk. Future applications will use long duration synthetic light curves that can span the dynamical time of the BLR. Using other empirical light curves, such as those from the Sloan Digital Sky Survey (SDSS) V Reverberation Mapping Campaign, can provide direct comparisons between the behaviors of our synthetic broad emission line profiles and those that are observed, while also allowing exploration of the model’s parameter space. With light curves expected from future telescopes and surveys such as the Long Synoptic Survey Telescope (Ivezić et al., 2019), we can utilize light curves with baselines as small as 200 days up to months to interpret observed line profiles. In a forthcoming paper, we will present the results of a parameter space exploration, as well as the consequences of different combinations of parameters for velocity jitter and breathing.